OpenStudy (sushi121212):

What is the neutron : proton ratio for the nucleus 206/82Pb? 1. 1 : 2.51 2. 2 : 1 3. None of these 4. 2.51 : 1 5. 0.51 : 1 6. 1.51 : 1

OpenStudy (cuanchi):

206/82Pb 82= number of protons 206= number of protons + numbers of neutrons numbers of neutrons =206 - number of protons numbers of neutrons =206 - 82 = 124 neutron : proton ratio 124:82 = ?????

OpenStudy (sushi121212):

Is it 6, 1.51 : 1?

OpenStudy (cuanchi):

yes

OpenStudy (sushi121212):

Okay, thank you!!